A vehicle tire, 42 inches diameter, is spinning at a rate of 545 rpm. How fast is the tire spinning?
1
Expert's answer
2020-10-15T03:13:39-0400
"v=\omega R=0.5\omega d\\v=\frac{2\pi}{60}545 (21\cdot 0.0254)=30.4\frac{m}{s}"